# Ultrafast Solvent Dynamics Drives the Formation of the Hydrated Electron in Photoexcited Water

### v2

@@ 3 Conclusions
We find that most excitations to the first band are localized on a single water molecule, but a significant fraction are delocalized over up to five molecules. This behavior, intrinsically linked to the asymmetry of the hydrogen-bond network, may explain the distinct photochemical outcomes observed at different energies within the first absorption band68, 69, 18. Our excited state molecular dynamics simulations reveal two distinct photophysical decay pathways: Hydrogen Atom Transfer (HAT) and Proton Coupled Electron Transfer (PCET). For the HAT mechanism, we predicted lifetimes and quantum yields are in excellent agreement with experimental data. Moreover, we observed the formation of the hydronium radical, a transient species previously proposed in theoretical studies but, to the best of our knowledge, not yet observed experimentally and reported here for the first time starting from photoexcited pure liquid water. This species has been suggested as a possible precursor to the hydrated electron74, 18. While our simulations, performed in the electronic excited state, can not exclude the possibility that it could release an electron in the ground electronic state, further studies are needed to confirm this pathway.
@@
In the PCET pathway, the hydrated electron emerges in the excited state. Our results further reveal that this localization is driven by ultrafast (<< 1 ps) collective rotational and translational motions of water molecules, particularly those participating in hydrogen bonding with the electron. In its early stages, the hydrated electron is highly delocalized. However, as localization proceeds, the gyration radius approaches values consistent with experimental measurements for the equilibrated hydrated electron in the electronic ground state. These findings indicate that a significant degree of electron localization together with collective waters rearrangements, occurs in the excited state before non-radiative decay. Moreover, we also observe the formation of hydroxyl radical–hydronium ion pairs, which was recently observed experimentally35. We investigated how their spatial separation influences both the gyration radius of the hydrated electron and the evolving energy gap during the excited-state dynamics. This correlation is particularly relevant in the early stages of the process, when diffusion is limited and the interactions between species are strongest.
@@
Finally, our work provides new insights into the broad fluorescence emission associated with the hydrated electron, as reported in experiments. We demonstrate a clear correlation between the emission energy and the degree of localization of the electron. This relationship offers a fresh perspective on the optical behavior of the hydrated electron. In recent studies80, 81, 82, fluorescence emission has been observed in different water solutions. Although different mechanisms have been proposed, the hydrated electron has not been considered as a contributing factor underlying these experimental observations.
@@
In conclusion, our theoretical studies here have provided fresh insights into the underlying mechanisms associated with one the most fundamental processes in the chemical physics and physical chemistry of water, namely how light interacts with liquid water and the subsequent photochemistry. Through a synergy of data-driven approaches, our approach lays the ground-work to study the birth of the excited-state hydrated electron in different bulk aqueous salt solutions80, 81, 82, at aqueous interfaces14, 83 and various biological contexts such as radiation-induced DNA damage4, 5. On the more methodological side, it is interesting to explore whether our simulations can be used to develop machine-learning approaches 84, 85, 86, 87, 88 to examine the excited-state dynamics of aqueous systems which would allow for expanding the sampling and statistics. This is an area we are actively pursuing and will be the subject of a future study.
@@ 4.1 Ground-State Molecular Dynamics
We began by constructing a periodic simulation box containing 64 water molecules with a side length of 12.42 Å, corresponding to a density of 1 $g/cm^{3}$ at room temperature. An initial equilibration was performed using an empirical water model, SPC/E89 with the GROMACS software90 within the NVT ensemble for 100 ns. The canonical-sampling velocity-rescaling91 thermostat was applied with a coupling time of 0.1 ps.
From this classical equilibrated trajectory, we selected 100 configurations, separated by 1 ns, as initial conditions to perform ab initio quality ground-state molecular dynamics simulations with a machine-learning interatomic potential (MLIP). Specifically, we chose an MLIP recently trained and developed using the DeePMD-kit architecture92 on data computed using Density Functional Theory (DFT) with the SCAN0 functional93, which incorporates $10\%$ exact exchange to mitigate the self-interaction error. Prior studies93 have shown that SCAN0 provides an excellent description of the structural, dynamical, and electronic properties (such as IR spectra, electronic band structure, and band gap) of bulk liquid water compared to the original SCAN meta-GGA functional94, 95. Each of the 100 simulations conducted with the MLIP potential was run for 1 ns to generate well-converged and decorrelated initial conditions (positions and velocities), leveraging DeePMD-kit92 in combination with the LAMMPS package96.
@@ 4.2 Excited-State Molecular Dynamics
The excited state molecular dynamics (ESMD) simulations were initiated from the first electronic excited state ($S_{1}$), computed using the Restricted Open-shell Kohn–Sham (ROKS) method 46, 47. ROKS has proven to be a reliable approach for describing dissociative and charge-transfer states in closed-shell systems 48, 49, 50, processes that play a central role in the photochemistry of water.
@@
To propagate the system on the $S_{1}$ potential energy surface, we employed the Atomic Simulation Environment (ASE)97, a python module which is interfaced with CP2K software98. At each timestep, we calculated the energy on the electronic ground state $S_{0}$ using Density Functional Theory (DFT) to account for possible non-radiative relaxation pathways51, 49 (see below) which in our current setup is determined by examining the energy gap. The simulations were performed in the ensemble NVE using a timestep of 0.5 fs and velocity verlet algorithm to perform the nuclei propagation of the system.
@@
Both excited and ground state calculations were performed using the hybrid PBEh(40)-rVV10 functional99 in combination with the DZVP-MOLOPT-SR basis set100. This level of electronic structure theory has been successfully applied to the study of liquid water101, 102, 103, demonstrating its accuracy in reproducing ionization potential, redox potentials, electronic levels of excess electron. Carter et. al51. also reported excellent agreement with experiments for aqueous iodine solutions using the same functional within the ROKS framework. More recently the functional PBEh(40) was identified as one of the most reliable choices for describing the hydrated electron in its electronic ground state71. In our current work, the favorable agreement with reproducing the experimental UV-absorption and other quantities sensitive to the excited states of liquid water gives us confidence in the quality of our results.
@@
Each ESMD simulation was propagated until the system reached a crossing point, defined as the step where the energy gap between the first excited state ($S_{1}$) and the ground state ($S_{0}$) dropped below 0.2 eV. This criterion provides a robust and efficient way to identify non-radiative decay events and allows us to characterize the distribution of deactivation pathways, offering a lower bound on the excited-state lifetimes64, 65. A known limitation of this approach is that it does not explicitly include non-adiabatic coupling vectors (NACVs), which are key quantities for describing electronic transitions during non-radiative decay. However, a previous study on the radiolysis of water cluster in vacuum showed that NACVs increase significantly as the energy gap decreases104. This strong correlation implies that small energy gaps are reliable indicators of high non-adiabatic transition probabilities. Therefore, by using the energy gap as a decay criterion, we effectively capture the relevant non-radiative dynamics, justifying the omission of explicitly determining NACVs in our simulations.
@@
The simulations that did not exhibit a decay were continued until a maximum time of 1 ps was reached. Spin density cube files were calculated and outputted every 5 fs during all ESMD trajectories and at the step of crossing, for post-processing analysis (see section below). All the electronic calculations in the present work were performed using CP2K version 2024.298.
@@ 5.1 Coordination Numbers
Here, $i$ is the atom index, while $\kappa=10$, $\textbf{r}_{c}=1.38$ Å are parameters taken from previous works used to study the dissociation of water in the ground-state to form hydronium and hydroxide ions105, 106, 107, 108. In this manner, for an isolated hydrogen atom not attached to any water molecule, the $CN_{H}$ is zero while for a hydrogen covalently bonded to a water molecule $CN_{H}$ is 1. In the case of $CN_{O}$, it takes on a value of 3, 2 and 1, for a hydronium (ion or radical), neutral molecular water and hydroxyl species respectively.
@@ 5.2 Hydrogen Bond (HB) Network Classification
HB were identified using a geometrical criteria introduced by Kumar et. al.109: O-O distance of $\leq 3.5$ Å and H-O—O angle $\leq 30\ deg$. A water molecule with four HB is classified as part of a normal HB network, otherwise, it is labeled as a defect (see Figure 1C for categories). Sensitivity analysis to the choice of these geometric parameters is addressed in Section S3 of the Supplementary Information showing that while the hydrogen-bond criteria naturally affect the exact populations of defects, the trends and physical insights inferred are robust.
@@ 5.3 Spin-Density Analysis for Excitation Localization
where $r_{0}^{i}$ is the position of the ithi-th oxygen atom and $R=1$ Å. To estimate the degree of localization, we computed the Inverse Participation Ratio (IPR)53, 52:
@@
This metric indicates how many water molecules participate in the excitation. The distribution of $IPR^{-1}$ values for all the initial configurations is shown in Figure 1B. The sensitivity analysis and validation to the choice of the integration radius parameter are discussed in section S2 in the Supplementary Information.
@@ 5.4 HAT versus PCET Mechanism
$H_{2}O\to HO^{\bullet}+H^{\bullet}\ \textbf{(HAT)}$ (4)
$H_{2}O\to HO^{\bullet}+H^{+}+e^{-}\ \textbf{(PCET)}$ (5)
@@
A direct calculation of the electron center from the spin density cube files yields a position located between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H}^{\bullet}$ (Equation 4), or between the $\mathrm{HO}^{\bullet}$ and the excess electron (Equation 5). Figure S5 in the Supplementary Information illustrates this point. This result lacks a clear physical interpretation, as it does not correspond to the actual location of any species and essentially implies that the spin density is located close to the geometric center of the species involved. This problem does not arise in earlier theoretical studies28, 25, 45, 73, as those simulations typically start with a pre-inserted electron in bulk water and the formation of an additional radical species are not observed.
@@
To overcome this challenge, we first identify the $\mathrm{HO}^{\bullet}$ species by locating the water molecule with the longest OH bond in each trajectory (see Section S6 of the Supplementary Information). Once identified, we remove the contribution of the $\mathrm{HO}^{\bullet}$ to the spin density by excluding a spherical region of radius 1.5 Å centered on its oxygen atom (see Section S7 in the Supplementary Information). This step yields a modified cube file that provides a more meaningful representation of the spin density to localize the hydrated excess electron. We then compute the electron center and gyration radius from this file using the protocol described in previous studies45 (see Section S8 in the Supplementary Information). Since the cube files were saved every 5 fs and at the crossing geometry, the electron center and gyration radius for the intermediate frames were obtained using linear interpolation between consecutive cube files.
@@
Once all the species were correctly identified, we classified each trajectory according to its reaction mechanism. For this, we analyzed the final frame of each simulation (corresponding to the geometry at the point of non-radiative crossing) and we computed the distance between the electron center and the nearest hydrogen atom. If the distance was less than $0.6$ Å, the trajectory was classified as hydrogen atom transfer (HAT), otherwise it was assigned to the proton coupled electron transfer (PCET). The $0.6$ Å threshold was chosen based on the radial distribution function (RDF) between the electron and all hydrogen atoms (see Figure S9 in the Supplementary Information). Specifically, it represents the midpoint between two cases: a bound hydrogen atom (distance electron and nearest H $\sim$0 Å) and the first maximum associated with hydrogen atoms in the solvation shell of the electron (distance between electron and nearest H $\sim$1.2 Å)45, 51. A short electron–hydrogen distance indicates that both species occupy the same spatial region, consistent with the formation of a neutral hydrogen atom. In contrast, larger distances suggest that the electron remains solvated while the proton is bound to nearby water molecules, consistent with a PCET process.
@@
G.D.M, D.D. and A.H. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101043272 – HyBOP). The views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. G.D.M., C.M. and A.H. also acknowledge MareNostrum5 (project EHPC-EXT-2023E01-029) for computational resources. C.J.M acknowledges support from the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Chemical Sciences, Geosciences, and Biosciences Division, Chemical Physics and Interfacial Sciences Program, FWP 16249.
@@
Supplementary Information for: Ultra-fast Solvent Dynamics Drives the Formation of the Hydrated Electron in Photoexcited Water
@@
Figure S1: Initial photo-absorption in liquid water: The histogram of energy gaps calculated with ROKS is shown as light-blue bars, while the histogram of energy gaps weighted by oscillator strength from TDDFT calculations is presented as orange bars. Both theoretical spectra were scaled (while preserving relative intensities) to better visibility and comparison with the experimental absorption spectrum, depicted as a red line. The experimental spectrum was digitalized from Ref 18.
@@
Using the modified cube files for the spin density (right panels in Figure S5) we can now determine the electron center and radius of gyration using the following expressions45:
@@
where $\rho^{s}$, $r_{c}$, $\lambda_{i}$ and $r_{g}$ represent the spin density, electron center, the eigenvalues of the second moment tensor S and the gyration radius, respectively. It is important to note that the previous definitions are not strictly valid under periodic boundary conditions45. To overcome this, we follow the approach proposed in previous studies45, in which the cube files are re-centered on the maximum of the spin density. This procedure ensures that the spin density decays to zero (or to very small values) at the edges of the simulation box, leading to more numerically stable results.
@@
Figure S9: Radial Distribution Function of the hydrated electron. Upper panel shows the the radial distribution function $g(\textbf{r})$ between the electron and the hydrogen atoms of the waters. Lower panel shows the $g(\textbf{r})$ for the electron and the oxygen atoms of the waters. The results obtained on the excited state with the ROKS method are shown in purple. The results using a recently reported machine-learning (ML) potential from Ref45 are shown in green. All the results for the excited state are for the case of low gyration radius (lower than 4 Å, see main text).

### v3

@@ 4 Conclusions
We find that most excitations to the first band are localized on a single water molecule, but a significant fraction are delocalized over up to five molecules. This behavior, intrinsically linked to the asymmetry of the hydrogen-bond network, may explain the distinct photochemical outcomes observed at different energies within the first absorption band92, 93, 117. Our excited state molecular dynamics simulations reveal two distinct photophysical decay pathways: Hydrogen Atom Transfer (HAT) and Proton Coupled Electron Transfer (PCET). For the HAT mechanism, we predicted lifetimes and quantum yields are in excellent agreement with experimental data. Moreover, we observed the formation of the hydronium radical, a transient species previously proposed in theoretical studies but, to the best of our knowledge, not yet observed experimentally and reported here for the first time starting from photoexcited pure liquid water. This species has been suggested as a possible precursor to the hydrated electron66, 117. While our simulations, performed in the electronic excited state, can not exclude the possibility that it could release an electron in the ground electronic state, further studies are needed to confirm this pathway.
@@
In the PCET pathway, the hydrated electron emerges in the excited state. Our results further reveal that this localization is driven by ultrafast (<< 1 ps) collective rotational and translational motions of water molecules, particularly those participating in hydrogen bonding with the electron. In its early stages, the hydrated electron is highly delocalized. However, as localization proceeds, the gyration radius approaches values consistent with experimental measurements for the equilibrated hydrated electron in the electronic ground state. These findings indicate that a significant degree of electron localization together with collective waters rearrangements, occurs in the excited state before non-radiative decay. Moreover, we also observe the formation of hydroxyl radical–hydronium ion pairs, which was recently observed experimentally62. We investigated how their spatial separation influences both the gyration radius of the hydrated electron and the evolving energy gap during the excited-state dynamics. This correlation is particularly relevant in the early stages of the process, when diffusion is limited and the interactions between species are strongest.
@@
Finally, our work provides new insights into the broad fluorescence emission associated with the hydrated electron, as reported in experiments. We demonstrate a clear correlation between the emission energy and the degree of localization of the electron. This relationship offers a fresh perspective on the optical behavior of the hydrated electron. In recent studies109, 110, 111, fluorescence emission has been observed in different water solutions. Although different mechanisms have been proposed, the hydrated electron has not been considered as a contributing factor underlying these experimental observations.
@@
In conclusion, our theoretical studies here have provided fresh insights into the underlying mechanisms associated with one the most fundamental processes in the chemical physics and physical chemistry of water, namely how light interacts with liquid water and the subsequent photochemistry. Through a synergy of data-driven approaches, our approach lays the ground-work to study the birth of the excited-state hydrated electron in different bulk aqueous salt solutions109, 110, 111, at aqueous interfaces108, 45 and various biological contexts such as radiation-induced DNA damage113, 11. On the more methodological side, it is interesting to explore whether our simulations can be used to develop machine-learning approaches 8, 6, 122, 7, 80 to examine the excited-state dynamics of aqueous systems. Such methods would enable simulations with larger unit cells and a greater number of trajectories, potentially improving agreement with experimental observations and revealing new physical insights.
@@ 5.1 Ground-State Molecular Dynamics
We began by constructing a periodic simulation box containing 64 water molecules with a side length of 12.42 Å, corresponding to a density of 1 $g/cm^{3}$ at room temperature. An initial equilibration was performed using an empirical water model, SPC/E9 with the GROMACS software1 within the NVT ensemble for 100 ns. The canonical-sampling velocity-rescaling13 thermostat was applied with a coupling time of 0.1 ps.
From this classical equilibrated trajectory, we selected 100 configurations, separated by 1 ns, as initial conditions to perform ab initio quality ground-state molecular dynamics simulations with a machine-learning interatomic potential (MLIP). Specifically, we chose an MLIP recently trained and developed using the DeePMD-kit architecture120 on data computed using Density Functional Theory (DFT) with the SCAN0 functional121, which incorporates $10\%$ exact exchange to mitigate the self-interaction error. Prior studies121 have shown that SCAN0 provides an excellent description of the structural, dynamical, and electronic properties (such as IR spectra, electronic band structure, and band gap) of bulk liquid water compared to the original SCAN meta-GGA functional19, 123. Each of the 100 simulations conducted with the MLIP potential was run for 1 ns to generate well-converged and decorrelated initial conditions (positions and velocities), leveraging DeePMD-kit120 in combination with the LAMMPS package103.
@@ 5.2 Excited-State Molecular Dynamics
The excited state molecular dynamics (ESMD) simulations were initiated from the first electronic excited state ($S_{1}$), computed using the Restricted Open-shell Kohn–Sham (ROKS) method 32, 77, 31. This method provides an alternative approach for calculating singlet excited states in closed-shell systems that are described by non-Aufbau electronic configurations. By minimizing the energy expression $E_{S}=2E_{mix}-E_{T}$, where $E_{S},E_{T},E_{mix}$ correspond to the energies of the singlet, triplet, and mixed states, respectively, the method effectively accounts for a two-determinant nature of the excited state. Consequently, ROKS offers a computationally efficient and reliable framework for describing dissociative and charge-transfer excited states50, 39, 30, 38.
@@
To propagate the system on the $S_{1}$ potential energy surface, we employed the Atomic Simulation Environment (ASE)59, a python module which is interfaced with CP2K software52. At each timestep, we calculated the energy on the electronic ground state $S_{0}$ using Density Functional Theory (DFT) to account for possible non-radiative relaxation pathways17, 39 (see below) which in our current setup is determined by examining the energy gap. The simulations were performed in the ensemble NVE using a timestep of 0.5 fs and velocity verlet algorithm to perform the nuclei propagation of the system.
@@
Both excited and ground state calculations were performed using the hybrid PBEh(40)-rVV10 functional112 in combination with the DZVP-MOLOPT-SR basis set107. This level of electronic structure theory has been successfully applied to the study of liquid water4, 2, 3, demonstrating its accuracy in reproducing ionization potential, redox potentials, electronic levels of excess electron. Excellent agreement with the experiments for aqueous iodine solutions were reported by Carter et. al.17 within ROKS framework and by Lan et. al.57 employing a similar approach with the same functional.
More recently the functional PBEh(40) was identified as one of the most reliable choices for describing the hydrated electron in its electronic ground state10. In our current work, the favorable agreement with reproducing the experimental UV-absorption and other quantities sensitive to the excited states of liquid water gives us confidence in the quality of our results.
@@
Each ESMD simulation was propagated until the system reached a crossing point, defined as the step where the energy gap between the first excited state ($S_{1}$) and the ground state ($S_{0}$) dropped below 0.2 eV. This criterion provides a robust and efficient way to identify non-radiative decay events and allows us to characterize the distribution of deactivation pathways, offering a lower bound on the excited-state lifetimes79, 43. A known limitation of this approach is that it does not explicitly include non-adiabatic coupling vectors (NACVs), which are key quantities for describing electronic transitions during non-radiative decay. However, a previous study on the radiolysis of water cluster in vacuum showed that NACVs increase significantly as the energy gap decreases96. This strong correlation implies that small energy gaps are reliable indicators of high non-adiabatic transition probabilities. Therefore, by using the energy gap as a decay criterion, we effectively capture the relevant non-radiative dynamics, justifying the omission of explicitly determining NACVs in our simulations23, 81. To further support this approximation, we evaluated the Landau-Zener (LZ) transition probabilities104, which in addition to the energy gap criterion, also account for the curvature of the potential energy surfaces (see section S10 in the Supplementary Information). This approach has been shown to reproduce the results consistent with those obtained using explicit NACVs in non-adiabatic dynamics97, 104, 116, 25, 68.
@@
The simulations that did not exhibit a decay were continued until a maximum time of 1 ps was reached. Spin density cube files were calculated and outputted every 5 fs during all ESMD trajectories and at the step of crossing, for post-processing analysis (see section below). All the electronic calculations in the present work were performed using CP2K version 2024.252.
@@ 6.1 Coordination Numbers
Here, $i$ is the atom index, while $\kappa=10$, $\textbf{r}_{c}=1.38$ Å are parameters taken from previous works used to study the dissociation of water in the ground-state to form hydronium and hydroxide ions94, 95, 91, 24. In this manner, for an isolated hydrogen atom not attached to any water molecule, the $CN_{H}$ is zero while for a hydrogen covalently bonded to a water molecule $CN_{H}$ is 1. In the case of $CN_{O}$, it takes on a value of 3, 2 and 1, for a hydronium (ion or radical), neutral molecular water and hydroxyl species respectively.
@@ 6.2 Hydrogen Bond (HB) Network Classification
HB were identified using a geometrical criteria introduced by Kumar et. al.55: O-O distance of $\leq 3.5$ Å and H-O—O angle $\leq 30\ deg$. A water molecule with four HB is classified as part of a normal HB network, otherwise, it is labeled as a defect (see Figure 1C for categories). Sensitivity analysis to the choice of these geometric parameters is addressed in Section S3 of the Supplementary Information showing that while the hydrogen-bond criteria naturally affect the exact populations of defects, the trends and physical insights inferred are robust.
@@ 6.3 Spin-Density Analysis for Excitation Localization
where $r_{0}^{i}$ is the position of the ithi-th oxygen atom and $R=1$ Å. To estimate the degree of localization, we computed the Inverse Participation Ratio (IPR)64, 114:
@@
This metric provides a continous measure of how many water molecules participate in the excitation. The distribution of $IPR^{-1}$ values for all the initial configurations is shown in Figure 1B. The sensitivity analysis and validation to the choice of the integration radius parameter are discussed in section S2 in the Supplementary Information.
@@ 6.4 HAT versus PCET Mechanism
$H_{2}O\xrightarrow{h\nu}HO^{\bullet}+H^{\bullet}\ \textbf{(HAT)}$ (4)
$H_{2}O\xrightarrow{h\nu}HO^{\bullet}+H^{+}+e^{-}\ \textbf{(PCET)}$ (5)
@@
A direct calculation of the electron center from the spin density cube files yields a position located between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H}^{\bullet}$ (Equation 4), or between the $\mathrm{HO}^{\bullet}$ and the excess electron (Equation 5). Figure S5 in the Supplementary Information illustrates this point. This result lacks a clear physical interpretation, as it does not correspond to the actual location of any species and essentially implies that the spin density is located close to the geometric center of the species involved. This problem does not arise in earlier theoretical studies42, 76, 58, 48, as those simulations typically start with a pre-inserted electron in bulk water and the formation of an additional radical species are not observed.
@@
To overcome this challenge, we first identify the $\mathrm{HO}^{\bullet}$ species by locating the water molecule with the longest OH bond in each trajectory (see Section S6 of the Supplementary Information). Once identified, we remove the contribution of the $\mathrm{HO}^{\bullet}$ to the spin density by excluding a spherical region of radius 1.5 Å centered on its oxygen atom (see Section S7 in the Supplementary Information). This step yields a modified cube file that provides a more meaningful representation of the spin density to localize the hydrated excess electron. We then compute the electron center and gyration radius from this file using the protocol described in previous studies58 (see Section S8 in the Supplementary Information). Since the cube files were saved every 5 fs and at the crossing geometry, the electron center and gyration radius for the intermediate frames were obtained using linear interpolation between consecutive cube files.
@@
Once all the species were correctly identified, we classified each trajectory according to its reaction mechanism. For this, we analyzed the final frame of each simulation (corresponding to the geometry at the point of non-radiative crossing) and we computed the distance between the electron center and the nearest hydrogen atom. If the distance was less than $0.6$ Å, the trajectory was classified as hydrogen atom transfer (HAT), otherwise it was assigned to the proton coupled electron transfer (PCET). The $0.6$ Å threshold was chosen based on the radial distribution function (RDF) between the electron and all hydrogen atoms (see Figure S9 in the Supplementary Information). Specifically, it represents the midpoint between two cases: a bound hydrogen atom (distance electron and nearest H $\sim$0 Å) and the first maximum associated with hydrogen atoms in the solvation shell of the electron (distance between electron and nearest H $\sim$1.2 Å)58, 17. A short electron–hydrogen distance indicates that both species occupy the same spatial region, consistent with the formation of a neutral hydrogen atom. In contrast, larger distances suggest that the electron remains solvated while the proton is bound to nearby water molecules, consistent with a PCET process.
Based on the preceding analysis, each trajectory can be unambiguously assigned to a specific mechanism. This allows us to determine the quantum yield and lifetime associated with each mechanism using the following expression:
$QY_{M}=\dfrac{\#\text{Trajectories with M}}{\#\text{Total trajectories}}*100$ (6)
$\tau_{M}=\dfrac{\sum_{i}^{N_{M}}t_{i}^{M}}{\#\text{Total trajectories with M}}$ (7)
where $M$ represent the mechanism (HAT or PCET), $QY,\tau$ are the quantum yields and lifetimes, $N_{M}$ is the number of trajectories with the $M$ mechanism and $t_{i}^{M}$ is the time at which the trajectory $i$ with the $M$ mechanism finds the ”crossing point”.
@@
G.D.M, D.D. and A.H. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101043272 – HyBOP). The views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. G.D.M., C.M. and A.H. also acknowledge MareNostrum5 (project EHPC-EXT-2023E01-029) for computational resources. C.J.M was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), the Chemical Sciences, Geosciences, and Biosciences Division, Chemical Physics and Interfacial Sciences Program, FWP 16249.
@@
Supplementary Information for: The Photochemical Birth of the Hydrated Electron in Liquid Water
@@
Figure S1: Initial photo-absorption in liquid water: The histogram of energy gaps calculated with ROKS is shown as light-blue bars, while the histogram of energy gaps weighted by oscillator strength from TDDFT calculations is presented as orange bars. Both theoretical spectra were scaled (while preserving relative intensities) to better visibility and comparison with the experimental absorption spectrum, depicted as a red line. The experimental spectrum was digitalized from Ref 117.
@@
Using the modified cube files for the spin density (right panels in Figure S5) we can now determine the electron center and radius of gyration using the following expressions58:
@@
where $\rho^{s}$, $r_{c}$, $\lambda_{i}$ and $r_{g}$ represent the spin density, electron center, the eigenvalues of the second moment tensor S and the gyration radius, respectively. It is important to note that the previous definitions are not strictly valid under periodic boundary conditions58. To overcome this, we follow the approach proposed in previous studies58, in which the cube files are re-centered on the maximum of the spin density. This procedure ensures that the spin density decays to zero (or to very small values) at the edges of the simulation box, leading to more numerically stable results.
@@
Figure S9: Radial Distribution Function of the hydrated electron. Upper panel shows the the radial distribution function $g(\textbf{r})$ between the electron and the hydrogen atoms of the waters. Lower panel shows the $g(\textbf{r})$ for the electron and the oxygen atoms of the waters. The results obtained on the excited state with the ROKS method are shown in purple. The results using a recently reported machine-learning (ML) potential from Ref58 are shown in green. All the results for the excited state are for the case of low gyration radius (lower than 4 Å, see main text).
S9 Validation of functional and basis set in ROKS approach
The main objective of this section is to assess the sensitivity of the ROKS approach to variations in the basis set and exchange–correlation functional. To this end, we selected two representative trajectories: one corresponding to the HAT mechanism and one to the PCET mechanism, originally computed using the double-zeta basis set (DZVP)107 and the PBEh(40)-rVV10 functional112 here referred to as PBEh (additional details of the calculation are provided in Computational Methods section). These trajectories were then recalculated using a triple-zeta basis set (TZV2P)107 and the CAM-B3LYP functional118. To reduce computational cost, the PCET trajectory was recomputed at 5 fs intervals, while the HAT trajectory was recalculated at every time step. The results are summarized in Figure S10.
Figure S10: Basis set and exchange-correlation functional validation within ROKS framework. Comparison between the DZVP (green line) and TZV2P (red line) basis sets using the PBEh(40)-rVV10 functional (PBEh) is shown in Panel A and Panel B for the HAT and PCET mechanisms, respectively. Similarly, comparison between the PBEh (green line) and CAM-B3LYP (CAM, red line) functionals using the DZVP basis set is shown in Panel C and Panel D for the HAT and PCET mechanisms, respectively. For the PCET trajectory, the PBEh-TZV2P (red line in Panel B) and CAM-DZVP (red line, Panel D) calculations were permorfed at configurations sampled every 5 fs to reduce the computational cost.
S10 Incorporating Non-Adiabatic Effects into the Photochemistry of Liquid Water
In this work, we employed an energy gap threshold to determine whether the system remains in the excited state or undergoes a non-radiative transition to the electronic ground state. This methodology is widely used in non-adiabatic dynamics when non-adiabatic coupling vectors (NACVs) are not available23, 81, as is the case in the present ROKS framework. To further assess the robustness of our approach, we additionally computed the Landau–Zener (LZ) probabilities associated with the non-radiative $S_{1}\to S_{0}$ transition. Unlike the simple energy gap criterion, the LZ approach incorporates both the energy gap and the curvature of the potential energy surfaces (see equation below):
$P_{ij}^{LZ}=exp\left(-\dfrac{\pi}{2\hbar}\sqrt{\dfrac{\Delta E_{ij}^{3}}{d^{2}\Delta E_{ij}/dt^{2}}}\right)$ (11)
where $\Delta E_{ij}$ is the energy gap between the states $i$ and $j$.
This method has been shown to closely reproduce results obtained from simulations that explicitly include NACVs across a wide range of systems and electronic structure levels97, 104, 116, 25, 68.
Figure S11 presents the evolution of the LZ probabilities and the energy gap for representative trajectories of the HAT (Panel A) and PCET (Panel B) mechanisms. In addition, Panel C shows a 2D density plot of the LZ probabilities versus the energy gap for all trajectories analyzed in this work.
Figure S11: Landau-Zener (LZ) Probabilities. Energy gap (purple line) and LZ probabilities (red line) as a function of time for representative HAT (Panel A) and PCET (Panel B) trajectories. Panel C: shows a 2D density plot of the LZ probabilities versus energy gap for all trajectories, with the inset providing a zoomed-in view at small LZ values. Red points indicate the values at the $S_{1}\to S_{0}$ crossing events.
As shown in Figure S11, the LZ probabilities increase sharply when the energy gap becomes small, confirming the consistency and validity of our energy gap based criterion for both mechanisms. For the ensemble of trajectories considered in this work, the majority of non-radiative transitions occur at high LZ probability values, as highlighted by the red points in Figure S11C.